OpenStudy (anonymous):

Solve. 8 = 26(x + 4)

OpenStudy (anonymous):

Just distribute 26 to the parenthesis first..

OpenStudy (anonymous):

26x+104?

OpenStudy (anonymous):

yes: 8 = 26x + 104 Now subtract 8 both the sides..

OpenStudy (shaik0124):

x= -96/26

OpenStudy (anonymous):

Do Subtract 8 from 26x or 108?

OpenStudy (shaik0124):

subtract from 8

OpenStudy (anonymous):

Like this: \(8 - 8 = 26x + 104 - 8\) Solve now..

OpenStudy (anonymous):

My final answer is -4?

OpenStudy (shaik0124):

final result is -96/26

OpenStudy (anonymous):

^^^ It can't be -96/26

OpenStudy (anonymous):

My choices are x = –4 x = –1 x = 0 x = 7

OpenStudy (anonymous):

Check your question once again.. 8 = 26(x+4) right ??

OpenStudy (shaik0124):

it is approx -4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{-48}{13} = -3.692 \approx -4\]

OpenStudy (shaik0124):

bcoz when u divide 96/26 u will get 3.69 so we can conclude it to 4

OpenStudy (anonymous):

I was right! Thanks guys!

OpenStudy (shaik0124):

pellet